OpenStudy (anonymous):

write an equation of the line through A (2, -3) and B (1,-1).

OpenStudy (lawliet):

\[(y+3)/(x-2)=(-1-(-3))/(1-2)=>y+3=-2(x-2)=>y+2x-1=0 \] is the equation

OpenStudy (anonymous):

find the slope, put it into point-slope form. Bingo. You get the equation!